# ECAFormer: Low-light Image Enhancement using Cross Attention

### v2

@@ Abstract
Low-light image enhancement (LLIE) is critical in computer vision. Existing LLIE methods often fail to discover the underlying relationships between different sub-components, causing the loss of complementary information between multiple modules and network layers, ultimately resulting in the loss of image details. To beat this shortage, we design a hierarchical mutual Enhancement via a Cross Attention transformer (ECAFormer), which introduces an architecture that enables concurrent propagation and interaction of multiple features. The model preserves detailed information by introducing a Dual Multi-head self-attention (DMSA), which leverages visual and semantic features across different scales, allowing them to guide and complement each other. Besides, a Cross-Scale DMSA block is introduced to capture the residual connection, integrating cross-layer information to further enhance image detail. Experimental results show that ECAFormer reaches competitive performance across multiple benchmarks, yielding nearly a 3% improvement in PSNR over the suboptimal method, demonstrating the effectiveness of information interaction in LLIE.
Introduction
@@
Capturing images in low-light conditions presents significant challenges in computer vision, including the degradation of fine details, reduced color saturation, diminished contrast and dynamic range, and uneven exposure. These issues can greatly compromise the integrity and clarity of captured data, negatively impacting the effectiveness of subsequent applications such as autonomous driving systems (Ranft and Stiller 2016; Cai et al. 2020; Bresson et al. 2017) and nighttime surveillance efforts (Fu et al. 2021). Naturally, enhancing the visibility of objects and details in low-light images is crucial, possessing far-reaching implications for a spectrum of vision applications.
@@
Currently, efforts for LLIE have predominantly utilized deep-learning techniques, which can be categorized into three main approaches: Conventional, Guided, and Two-Branch Methods, as depicted in Figure 1. Conventional methods, such as LLNet (Lore, Akintayo, and Sarkar 2017) and UFormer (Wang et al. 2022), hinge on data-driven strategies and tailored loss functions but are constrained by data complexity, hindering the model’s ability to capture robust representations. To overcome this, guided methods integrate prior knowledge into LLIE. They decompose the input into essential and target components, leveraging the former knowledge to guide the enhancement of the latter through specialized networks. For example, some guided methods based on Retinex theory (Land and McCann 1971) focus on preserving the consistency of the reflection component while adjusting the illumination. Another line to integrate prior knowledge is the Two-Branch Methods (Ma et al. 2021; Tang et al. 2022), which further refine the input by separating it into distinct sub-components, each processed by dedicated enhancement networks. However, despite the use of prior knowledge, these methods may overlook the interrelationships among components in the latent space. This limitation, inherent in the model architecture, can result in a loss of image details during enhancement.
@@
Regarding the current situation in the LLIE field, we focus on enhancing the information interactions across different components and layers to fill in the image details. Specifically, we proposed a Hierarchical Mutual Enhancement via a Cross Attention transformer (ECAFormer) as shown in Figure 2. To achieve this, we employ the U-shaped architecture as the foundation of our model to enhance low-light images. Following the idea of two branch methods, we incorporate a knowledge-driven strategy that decomposes images into distinct visual and semantic components, which facilitates detail preserving with the guidance of semantic information. Further, to fully leverage the advantages of both, we integrate a modified transformer to harness visual and semantic components. In order to enhance the interactions across different sub-components in the latent space, we designed Dual Multi-head Self-Attention (DMSA) to handle multiple inputs, capturing potential relationships between different components. Notably, the transition from Multi-Head Self-Attention (MHSA) in the vanilla transformer to DMSA modules is strategically designed to simultaneously exploit both long- and short-range dependencies within the image data. Moreover, to further preserve the details, we also design a Cross-Scale DMSA to foster an interplay between residual and current layer information, which is ingeniously leveraged to refine the residual connections within ECAFormer. Our main contribution can be summarized as follows:
@@
  • We achieve competitive results on six LLIE datasets emphasizing computational frugality and parameter parsimony.
Related Work
LLIE: Low-light Image Enhancement
@@
Currently, the rapid advancement in low-light image dataset collection has led to the emergence of numerous deep learning-based enhancement methods (Li et al. 2021). LLNET (Lore, Akintayo, and Sarkar 2017) introduced a variant of the stacked sparse denoising autoencoder for improving degraded images, establishing a foundational framework for the application of deep learning in image enhancement. RetinexNet (Wei et al. 2018) employed a deep Retinex-based architecture to enhance low-light images by decomposing them into illumination and reflectance components. RUAS (Liu et al. 2021) utilized an advanced search unfolding technique based on a Retinex architecture. EnlightenGAN (Jiang et al. 2021) innovatively used a generative inverse network as the primary framework, initially training with unpaired images. LEDNet (Zhou, Li, and Loy 2022), is a robust network specifically designed to address the dual challenges of low-light enhancement and deblurring simultaneously.
However, these methods primarily rely on Convolutional Neural Networks (CNN) models as their backbone and do not fully leverage global information or feature interaction for LLIE, resulting in detail loss. This paper proposes a transformer model with cross attention module that utilizes the global perception ability of transformers to deeply fuse visual and semantic features.
Components Decomposition and Fusion
@@
Compared to conventional LLIE methods, the components decomposition and fusion techniques have more effectively captured the latent knowledge within the data and significantly enhanced model performance. Among these approaches, feature combination frameworks can be classified into two types, as illustrated in Figure 1: (1) Two-Branch Methods, which optimize distinct features using separate networks, such as CSDGAN (Ma et al. 2021) and DRLIE (Tang et al. 2022); (2) Guided Methods, which use one feature to guide the optimization of another feature, exemplified by SNRNet (Xu et al. 2022).
@@
However, the aforementioned model architectures fail to fully leverage the complementary information between the decomposed sub-components, often treating them too independently and hindering the network from learning the intrinsic relationships between multiple sub-components. Therefore, we propose the DMSA module, which can simultaneously handle and inter-relate two features during forward computation. This modification enables the model to learn the relationships between sub-components, integrating their respective advantages and mitigating biases from ideal assumptions, thus improving the final output of the model.
Figure 2: The flowchart of ECAformer
Methodology
@@
In this section, we present a comprehensive description of each network module proposed and the corresponding loss functions employed. Inspired by the successful U-shaped architecture in LLIE (Li et al. 2021), we introduce a transformer network with a U-shaped architecture, depicted in Figure 2. Specifically, ECAFormer comprises three main components: (1) Visual-semantic convolution module Visual-Semantic Convolution Module, which generates short-range features (visual-features) and long-range features (semantic-features). (2) The U-shaped cross attention Transformer DMSA interacts with long- and short-term input features through DMSA, simultaneously propagating these features. (3) Mapping convolution, where the module projects the interacted features back to image features ${\in\mathbb{R}^{\emph{C}\times\emph{H}\times\emph{W}}}$. In practice, the low-light input image undergoes convolution filtering to extract semantic information and visual detail information respectively. The enhanced image is then obtained through interaction enhancement of a U-shaped network using semantic and visual features.
Problem Definition
@@
Low-light image enhancement aims to mitigate the challenges associated with low brightness, low contrast, noise, and artifacts in images captured under inadequate lighting conditions. These issues arise from the loss of visual information due to insufficient illumination or complex environmental factors at the time of image capture. The objective of low-light image enhancement is to utilize models to restore and enhance low-light images $\mathbb{L}$, bringing them closer to normal images $\mathbb{N}$ and rendering them more perceptible to the human visual system.
@@
where ${I\in\mathbb{R}^{C\times H\times W}}$ represents the input low-light images $\mathbb{L}$, ${\hat{I}\in\mathbb{R}^{C\times H\times W}}$ represents the enhanced images, and ${\theta}$ represents the parameters of the network function $f$.
The objective of the optimization process is to find the optimal parameters ${\theta}$ that minimize the differences between the enhanced images and the reference high-light or normal images $\mathbb{N}$. This can be formulated as:
@@
Model Framework
Preliminary
@@
The attention mechanism facilitates interactivity among elements, significantly enhancing global feature extraction capabilities. The conventional self-attention mechanism is defined by Eq. 4, where the vectors $Q$, $K$, and $V$ represent query, key, and value derived from a single input. This operation selectively focuses on crucial information, optimizing resource utilization and swiftly capturing the most relevant data. Leveraging this advantage, we have developed a unique Dual Multi-head Self-Attention module (DMSA) to facilitate the fusion of two distinct features. This module processes two inputs and enables their interaction via the attention mechanism while preserving their dimensional integrity. A detailed discussion of this module will be provided later.
@@
We have developed a U-shaped network structure with enhanced residual information, specifically tailored to improve the network’s capacity for multi-scale interaction. This architecture facilitates a comprehensive synthesis of features across various scales. During the down-sampling phase, we strategically employed a 2$\times$ scaling factor to compel the network to undergo a rigorous compression process, as we separate the image into visual and semantic features. This intentional compression is crucial as it allows the network to extract more refined global information, which is essential for understanding broader contextual cues. The methodology and details of the down-sampling stage are illustrated in Eq. 5.
@@
where $\{i\in{1,2\}}$ denotes the step of the down-sample stage, $RS(\cdot)$ represents the resample operation, the $DMSA_{i}$ denotes the DMSA Block varies at $i$ stages. At the bottom of the U-shaped network, we used a bottleneck consisting of two DMSA modules. And in the up-sampling process, we employed operations proportional to down-sampling. During the up-sampling phase, we utilized CSDMSA to effectively retain and restore intricate details that are often lost during the down-sampling process through residual information. The residual information initially engages in cross attention with their corresponding features, followed by cross attention between the two types of features in the CSDMSA block. This process is crucial in maintaining the fidelity of feature representations, ensuring that the reconstructed outputs closely resemble the original inputs. Finally, we obtain the final output through concatenation and mapping convolution. It can be represented by Eq. 6.
@@
Visual-Semantic Convolution Module
@@
DMSA
Figure 4: Archtecture of DMSA
Self-attention is highly efficient at directing concentrated attention toward pivotal information, thereby optimizing resource utilization and rapidly acquiring the most relevant data. However, its standard configuration is restricted to processing only a single input, which can limit its applicability in scenarios requiring complex interactions between multiple data streams. To address this constraint, we have innovated the Dual Multi-head Self Attention (DMSA) module as shown in Figure 4, a sophisticated mechanism designed to effectively manage and integrate dual inputs. In this module, inputs $\alpha$ and $\beta$ are processed to generate distinct sets of $Q$, $K$, and $V$ vectors through separate mappings. We strategically cross the key vectors of both inputs within this module to enhance their inter-connectivity, fostering a richer, more comprehensive interaction. This is achieved by the multiplication of the query vector with the corresponding crossed key vector, producing an attention map that encapsulates the intricate dynamics between the two inputs. Enhanced by the adaptive sampling operations of the U-shaped network, this setup facilitates a profound interaction across multiple scales, allowing for a nuanced integration of features that is essential for complex analysis and interpretation tasks.
@@
The scaling factor ${\zeta}$ is determined through optimization. To preserve the spatial positional relationships among pixels, we have incorporated a position embedding ($PosEmb$) module, executed via convolution operations. Consequently, the output of the Dual Multi-head Self Attention module $[\alpha^{\prime},\beta^{\prime}]$ can be formally defined in Eq. 8.
@@
Loss Function
We incorporated two types of loss functions that better align with human visual perception and facilitate faster model training. The Total Loss $\mathcal{L}_{Total}$ can be represented as Eq. 12, $\lambda\in{[0,1]}$.
@@
Perceptual Loss
Perceptual Loss (Johnson, Alahi, and Fei-Fei 2016) adopts an efficient approach by quantifying the discrepancies through the squared error between features extracted from specific layers or an aggregation of multiple layers after both the ground truth and the reconstructed image have traversed the same pre-trained neural network. The Perceptual Loss $\mathcal{L}_{p}$ is given as follows:
@@
Benefiting from the addition of epsilon, the phenomenon of gradient vanishing when $\hat{I}$ and R are very close has been alleviated, making the model easier to train.
@@
UF RN EnGAN RUAS DRBN KinD RSTM LLF MIRNet SNR ECAFormer
LOL-v1 16.36 16.77 17.48 18.23 20.13 20.86 22.43 23.65 24.14 24.61 24.24
LOL-v2-r 18.82 15.47 18.23 18.37 20.29 14.74 19.94 20.06 20.02 21.48 21.99
LOL-v2-s 19.66 17.13 16.57 16.55 23.22 13.29 21.41 24.04 21.94 24.14 25.86
SID 18.54 16.48 17.23 18.44 19.02 18.02 22.27 - 20.84 22.87 24.69
SMID 27.20 22.83 22.62 25.88 26.6 22.18 26.97 - 25.66 28.49 29.34
SDSD-in 23.17 20.84 20.02 23.17 24.08 21.95 25.67 - 24.38 29.44 29.11
SDSD-out 23.85 20.96 20.10 23.84 25.77 21.97 24.79 - 27.13 28.66 29.57
Average 21.08 18.64 18.89 20.64 22.73 19.00 23.35 - 23.44 25.67 26.40
FLOPS(G) 12.00 587.47 61.01 0.83 48.61 34.99 144.25 22.52 785.00 26.35 27.34
Params(M) 5.29 0.84 114.35 0.003 5.27 8.02 26.13 24.55 31.76 4.01 2.50
Table 1: Comparison of PSNR across 7 public datasets using different methods. ECAFormer achieved promising results with the least number of parameters.
Experiment

### v3

# ECAFormer: Low-light Image Enhancement using Dual Cross Attention

@@ Abstract
Low-light image enhancement (LLIE) aims to improve the perceptibility and interpretability of images captured in poorly illuminated environments. Existing LLIE methods often fail to capture the non-local self-similarity and long-range dependencies, causing the loss of complementary information between multiple modules or network layers, ultimately resulting in the loss of image details. To alleviate this issue, we design a hierarchical mutual Enhancement via a dual cross-attention transformer (ECAFormer), which introduces an architecture that enables concurrent propagation and interaction of multiple disentangling features. To capture the non-local self-similarity, we design a Dual Multi-head self-attention (DMSA), which leverages the disentangled visual and semantic features across different scales, allowing them to guide and complement each other. Further, a cross-scale DMSA block is incorporated to capture residual connections, thereby integrating cross-layer information and capturing the long-range dependencies. Experimental results show that the ECAFormer reaches competitive performance across multiple benchmarks, yielding nearly a 3.7% improvement in PSNR over the suboptimal method, demonstrating the effectiveness of information interaction in LLIE. For facilitating the efforts to replicate our results, our implementation is available on GitHub  [^1] https://github.com/ruanyudi/ECAFormer..
keywords:
Low-light image enhancement, Cross Attention, Transformer.
[…]
1 Introduction
@@
Capturing images in low-light conditions presents significant challenges in computer vision, including the degradation of fine details, reduced color saturation, diminished contrast and dynamic range, and uneven exposure. These issues can greatly compromise the integrity and clarity of captured data, negatively impacting the effectiveness of subsequent applications such as autonomous driving systems [1, 2, 3] and nighttime surveillance efforts [4, 5]. Naturally, enhancing the visibility of objects and details in low-light images is crucial, possessing far-reaching implications for a spectrum of vision applications.
@@
Currently, efforts for LLIE have predominantly utilized deep-learning techniques, which can be categorized into three main approaches: Conventional, Guided, and Two-Branch Methods, as depicted in Figure 1. Conventional methods, such as LLNet [6] and UFormer [7], hinge on data-driven strategies and tailored loss functions but are constrained by data complexity, hindering the model’s ability to capture robust representations. To overcome this, guided methods [8, 9] integrate prior knowledge into LLIE. They decompose the input into essential and target components, leveraging the former knowledge to guide the enhancement of the latter through specialized networks. For example, some guided methods [10, 11] based on Retinex theory [12] focus on preserving the consistency of the reflection component while adjusting the illumination. Another method for incorporating prior knowledge is the Two-Branch Methods, which involve refining the input by dividing it into separate sub-components and processing each component with its dedicated enhancement network [13, 14, 15, 16]. However, despite the use of prior knowledge, these methods fail to capture the non-local self-similarity and long-range dependencies. This limitation, inherent in the model architecture, can result in a loss of image details during enhancement.
@@
Regarding the current situation in the LLIE field, we focus on enhancing the information interactions across different disentangled components and layers to fill in the image details. To achieve this, we employ the U-shaped architecture as the foundation of our model to enhance low-light images. Specifically, we proposed ECAFormer as shown in Figure 2. Following the idea of two branch methods, we incorporate a knowledge-driven strategy that disentangles the images into distinct visual and semantic components, which facilitates detail preserving with the guidance of semantic information. To capture non-local similarities, we emphasize leveraging the interactive advantages of both visual and semantic components by integrating a modified transformer that effectively harnesses these elements. Specifically, to enhance the interactions across different sub-components in the latent space, we designed DMSA to handle multiple inputs, which captured potential relationships between different components. Notably, the transition from Multi-Head Self-Attention (MHSA) in the vanilla transformer to DMSA modules is strategically designed to simultaneously exploit both long- and short-range dependencies within the image data. Furthermore, to capture long-range dependencies, we designed a Cross-Scale DMSA that facilitates the interaction between residual and current layer information. This mechanism is ingeniously leveraged to refine the residual connections within ECAFormer. Our main contribution can be summarized as follows:
  • 1.
@@
    We achieve competitive results on six LLIE datasets, emphasizing computational frugality and parameter parsimony.
The paper is structured as follows. In Section 2, we respectively reviewed the related work in the field of low light enhancement and cross-attention mechanisms. In Section 3, we introduced the problem definition, the proposed method, and the details of model optimization. In Section 4, we compared our model with other methods using six publicly available datasets and conducted ablation experiments to validate its effectiveness. Finally, we summarized in Section 5 and proposed directions for future work.
2 Related Work
2.1 Convolutional Neural Networks based LLIE
@@
The rapid advancement in low-light image dataset collection has led to the emergence of numerous deep learning-based enhancement methods [17, 18]. LLNET [6] introduced a variant of the stacked sparse denoising autoencoder for improving degraded images, establishing a foundational framework for the application of deep learning in image enhancement. RetinexNet [19] employed a deep Retinex-based architecture to enhance low-light images by decomposing them into illumination and reflectance components. RUAS [20] utilized an advanced search unfolding technique based on a Retinex architecture. EnlightenGAN [21] innovatively used a generative inverse network as the primary framework, initially training with unpaired images. LEDNet [22], is a robust network specifically designed to address the dual challenges of low-light enhancement and deblurring simultaneously. LACN [23] proposed a novel module by introducing the attention mechanism into the ConvNeXt backbone network.
However, these Convolutional Neural Networks (CNN) based models fail to capture long-range dependencies, resulting in the loss of detailed information. To alleviate this, this paper introduces the transformer architecture and specifically designed cross-scale DMSA block, which significantly enhances the model’s ability to extract long-range dependencies and global information.
2.2 Disentangling based LLIE
@@
Compared to conventional LLIE methods [24, 25, 26], the disentangling-based techniques have more effectively captured the latent knowledge within the data and significantly enhanced model performance. Among these approaches, it can be classified into two types, as illustrated in Figure 1: (1) Two-Branch Methods, which optimize distinct features using separate networks, such as CSDGAN [13] and DRLIE [14]; (2) Guided Methods, which use one feature to guide the optimization of another feature, such as SNRNet [27] and RetinexFormer [28].
@@
However, the aforementioned model architectures inadequately exploit the complementary information among the decomposed sub-components, frequently addressing them in isolation. This approach hinders the network’s ability to learn the non-local self-similarity and the intrinsic relationships between multiple sub-components. Therefore, we propose the DMSA module, which can simultaneously handle and inter-relate two features during forward computation. This modification enables the model to learn the relationships between sub-components, integrating their respective advantages, thus effectively capturing the non-local self-similarity and the intrinsic relationships to improve the final output of the model.
@@
Figure 2: The flowchart of ECAformer mainly consists of three parts: (1) Visual-Semantic Convolution modules 3.2.2, which output short-range features (visual-features) and long-range features (semantic-features). (2)The U-shaped cross-attention Transformer 3.2.3 engages with long- and short-term input features through DMSA, concurrently propagating these features. (3) Mapping convolution, where the module projects the interacted features back to image features.${\in\mathbb{R}^{\emph{C}\times\emph{H}\times\emph{W}}}$.
3 Methodology
In this section, we present a comprehensive description of each network module proposed and the corresponding loss functions employed. Inspired by the successful U-shaped architecture in LLIE [17], we introduce a transformer network with a U-shaped architecture, depicted in Figure 2. Specifically, ECAFormer comprises three main components: (1) Visual-semantic convolution module 3.2.2, which generates short-range features and long-range features. (2) The U-shaped cross attention Transformer 3.2.3 interacts with long- and short-term input features through DMSA, simultaneously propagating these features. (3) Mapping convolution, where the module projects the interacted features back to image features ${\in\mathbb{R}^{\emph{C}\times\emph{H}\times\emph{W}}}$. In practice, the low-light input image undergoes convolution filtering to extract semantic information and visual detail information respectively. The enhanced image is then obtained through interaction enhancement of a U-shaped network using semantic and visual features.
3.1 Problem Definition
@@
Low-light image enhancement aims to mitigate the challenges associated with low brightness, low contrast, noise, and artifacts in images captured under inadequate lighting conditions. These issues arise from the loss of visual information under insufficient illumination or complex environmental factors at the time of image capture. The objective of low-light image enhancement is to utilize models to restore and enhance low-light images $\mathbb{L}$, bringing them closer to normal images $\mathbb{N}$ and rendering them more perceptible to the human visual system.
@@
where ${I\in\mathbb{R}^{C\times H\times W}}$ represents the input low-light images, ${\hat{I}\in\mathbb{R}^{C\times H\times W}}$ represents the enhanced images, and ${\theta}$ represents the parameters of the network function $f$.
The optimization objective is to find the optimal parameters ${\theta}$ that minimize the differences between the enhanced images and the reference high-light or normal images. This can be formulated as:
@@
3.2 Model Framework
3.2.1 Preliminary
@@
The attention mechanism facilitates interactivity among elements, significantly enhancing global feature extraction capabilities. The conventional self-attention mechanism is defined by Eq. (4), where the vectors $Q$, $K$, and $V$ represent query, key, and value derived from a single input. This operation selectively focuses on crucial information, optimizing resource utilization and swiftly capturing the most relevant data. Leveraging this advantage, we have developed a unique DMSA to facilitate the fusion of two distinct features. This module processes two inputs and enables their interaction via the attention mechanism while preserving their dimensional integrity. A detailed discussion of this module will be provided later.
@@
We have developed a U-shaped network structure with enhanced residual information, specifically tailored to improve the network’s capacity for multi-scale interaction. This architecture facilitates a comprehensive synthesis of features across various scales. During the down-sampling phase, we strategically employed a 2$\times$ scaling factor to compel the network to undergo a rigorous compression process, as we separate the image into visual and semantic features. This intentional compression is crucial as it allows the network to extract more refined global information, which is essential for understanding broader contextual cues. The methodology and details of the down-sampling stage are illustrated in Eq. (5).
@@
where $\{i\in{1,2\}}$ denotes the step of the down-sample stage, $RS(\cdot)$ represents the resample operation, the $DMSA_{i}$ denotes the DMSA Block varies at $i$ stages. At the bottom of the U-shaped network, we used a bottleneck consisting of two DMSA modules. And in the up-sampling process, we employed operations proportional to down-sampling. During the up-sampling phase, we utilized CSDMSA to effectively retain and restore intricate details that are often lost during the down-sampling process through residual information. The residual information initially engages in cross attention with their corresponding features, followed by cross attention between the two types of features in the CSDMSA block. This process is crucial in maintaining the fidelity of feature representations, ensuring that the reconstructed outputs closely resemble the original inputs. Finally, we obtain the final output through concatenation and mapping convolution. It can be represented by Eq. (6).
@@
3.2.2 Visual-Semantic Convolution Module
@@
3.2.3 DMSA
Figure 4: DMSA: A highly symmetrical module is depicted in the figure, which illustrates the process of computing the $[\alpha^{\prime},\beta^{\prime}]$ guided by each other using cross attention.
Self-attention is highly efficient at directing concentrated attention toward pivotal information, thereby optimizing resource utilization and rapidly acquiring the most relevant data. However, its standard configuration is restricted to processing only a single input, which can limit its applicability in scenarios requiring complex interactions between multiple data streams. To address this constraint, we have innovated the DMSA module as shown in Figure 4, a sophisticated mechanism designed to effectively manage and integrate dual inputs. In this module, inputs $\alpha$ and $\beta$ are processed to generate distinct sets of $Q$, $K$, and $V$ vectors through separate mappings. We strategically cross the key vectors of both inputs within this module to enhance their inter-connectivity, fostering a richer, more comprehensive interaction. This is achieved by the multiplication of the query vector with the corresponding crossed key vector as in Eq. (7), producing an attention map that encapsulates the intricate dynamics between the two inputs. Enhanced by the adaptive sampling operations of the U-shaped network, this setup facilitates a profound interaction across multiple scales, allowing for a nuanced integration of features that is essential for analysis and interpretation tasks.
@@ 3.2.3 DMSA
The scaling factor ${\zeta}$ is determined through optimization. To preserve the spatial positional relationships among pixels, we have incorporated a position embedding ($PosEmb$) module, executed via convolution operations. Consequently, the output of the Dual Multi-head Self Attention module $[\alpha^{\prime},\beta^{\prime}]$ can be formally defined in Eq. (8).
@@ 3.2.4 CSDMSA
3.3 Loss Function
We incorporated two types of loss functions that better align with human visual perception and facilitate faster model training. The Total Loss $\mathcal{L}_{Total}$ can be represented as Eq. (12), $\lambda\in{[0,1]}$.
@@ 3.3 Loss Function
3.3.1 Perceptual Loss
Perceptual Loss [29] adopts an efficient approach by quantifying the discrepancies through the squared error between features extracted from specific layers or an aggregation of multiple layers after both the ground truth and the reconstructed image have traversed the same pre-trained neural network. The Perceptual Loss $\mathcal{L}_{p}$ is given as follows:
@@ 3.3.2 Charbonnier Loss
Benefiting from the addition of $\epsilon$, the phenomenon of gradient vanishing when $\hat{I}$ and R are very close has been alleviated, making the model easier to train.
@@
Table 1: Comparison of PSNR of different methods across 5 public datasets. The data was obtained either by training with publicly available code or from the information provided in the paper. ECAFormer achieved promising results without a significant increase in parameters.
Methods LOL-v1 LOLv2-r LOLv2-s SID SMID AVG Params
RN(BMVC 2018) 16.77 15.47 17.13 16.48 22.83 17.74 0.84
KinD(ACMMM 2019) 20.86 14.74 13.29 18.02 22.18 17.82 8.02
EnGAN(TIP 2021) 17.48 18.23 16.57 17.23 22.62 18.43 114.35
RUAS(CVPR 2021) 18.23 18.37 16.55 18.44 25.88 19.49 0.00
DRBN(TIP 2021) 20.13 20.29 23.22 19.02 26.60 21.85 5.27
UF(CVPR 2022) 16.36 18.82 19.66 18.54 27.20 20.12 5.29
RSTM(CVPR 2022) 22.43 19.94 21.41 22.27 26.97 22.60 26.13
MIRNet(TPAMI 2022) 24.14 20.02 21.94 20.84 25.66 22.52 31.76
SNR(CVPR 2022) 24.61 21.48 24.14 22.87 28.49 24.32 4.01
LLF(AAAI 2023) 23.65 20.06 24.04 \ \ \ 24.55
RF(ICCV 2023) 23.69 21.43 24.16 23.85 27.55 24.14 1.61
U2E-Net(PR 2024) 23.92 \ \ \ \ \ \
ECAFormer 24.24 21.99 25.86 24.69 29.34 25.22 2.50
4 Experiment